Scores on a chemistry final exam are normally distributed with a mean of 280 and a standard deviation of 20. Determine the percentage of samples of size 4 that will have mean scores within 14 points of the population mean score of 280. Round to two decimal places.

A.

99.48 %

B.

83.84 %

C.

91.92 %

D.

51.60 %
Answer #1

Solution :

Given that,

mean = = 280

standard deviation = = 20

= / n = 20/ 4 = 10

= P[(-14) /10 < ( - ) / < (14) / 10)]

= P(-1.4 < Z < 1.4)

= P(Z < 1.4) - P(Z < -1.4)

= 0.9192 - 0.0808

= 0.8384 = 83.84%

percentage = 83.84%

option B. is correct